Grade 11Physical Chemistry

CN- in 200gms of KCN
calculate the moles???????????????????

harshat
10 Years agoGrade 11

1 Answer

Vikas TU
10 Years ago
moles of KCN = 200/65.12 = 3.07125 moles
moles of CN- = 3.07125*1 = 3.07125 are the moles of CN- in KCN.